OpenStudy (anonymous):

Let f and g be differentiable functions such that: f(1) = 4, g(1) = 3, f′(3) = -5, f′(1) = -4, g′(1) = -3, g′(3) = 2. If h(x) = f(g(x)), then h′(1)= ?

OpenStudy (anonymous):

use the chain rule, that \[(f(g(x))'=f'(g(x))\times g'(x)\] so \[(f(g(1))'=f'(g(1))\times g'(1)\] substitute the numbers and see what you get

OpenStudy (anonymous):

Is it (-4)(3) * (2)? Equalling -24?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it is not a product in the first part , it is \[f'(g(1))\]

OpenStudy (anonymous):

\[g(1)=3, f'(g(1))=f'(3)=-5\] so it is \[f'(g(1))\times g'(1)=-5\times (-3)

OpenStudy (anonymous):

\[f'(g(1))\times g'(1)=-5\times (-3)\]

OpenStudy (anonymous):

OH okay. That makes sense then with why they give you f′(3). Thank you.

OpenStudy (anonymous):

they gave you lots of extra information to see if you knew what to do